OpenStudy (anonymous):

write the slope-intercept inequality of the line!!! (6,3) (0,-1)

OpenStudy (anonymous):



OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

im confused...............the slope is -2/6 right?

OpenStudy (anonymous):

(6,3) (0,-1) need the slope, it is \[\frac{3-(-1)}{6-0}=\frac{4}{6}=\frac{2}{3}\] then you know the y intercept is \((0,-1)\) so you can put in in slope intercept form right away \[y=\frac{2}{3}x-1\]

OpenStudy (anonymous):

slope is positive

OpenStudy (anonymous):

you have \((0,-1)\) and \((6,3)\) so 6 units right (from 0 to 6) and 4 units UP from -1 to 3

OpenStudy (anonymous):

isnt it 1-3/0-6?

OpenStudy (anonymous):

or use \[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

slope is -4/-6

OpenStudy (anonymous):

you can do it that way, but your first coordinate is -1 not 1 right?

OpenStudy (anonymous):

so confusing!!!

OpenStudy (anonymous):

\[\frac{-1-3}{0-6}=\frac{-4}{-6}=\frac{2}{3}\]

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

ok so if the slope is 2/3

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

then what do i do after?

OpenStudy (anonymous):

y=2/3x+b how do i get b?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

this line you wrote isnt it 1-3/0-6? is wrong it should be -1-3/0-6

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

you are given b in the question. b is the y intercept, which is the second coordinate when \(x=0\)

OpenStudy (anonymous):

oh wait so

OpenStudy (anonymous):

y=2/3x+1?

OpenStudy (anonymous):

since you were given one point is \((0,-1)\) you know \(b=-1\)